# On the Expressiveness of Visual Prompt Experts

### v5

@@ 4.2 Adaptive Prompt
Efficiency Considerations. Following VPT, we insert adaptive prompts into every ViT block. While the feature projector is a lightweight MLP, distinct projector for each block would incur substantial parameter and memory overhead. To mitigate this, we employ a single, shared feature projector $g(\cdot)$ reused across all ViT blocks. This significantly reduces the overall computational cost and enhances efficiency, while preserving prompt expert expressiveness relative to VPT. Furthermore, an independent LayerNorm lei2016layer is incorporated, we incorporate an independent LayerNorm lei2016layer before the token-wise projectors in each block. Regarding parameter complexity, VPT introduces approximately $L\times N_{p}\times d$ parameters across $L$ ViT blocks. In contrast, VAPT’s learnable parameters include token-wise projectors, a channel-wise convolution layer per block, and a shared feature projector. The total parameter count is given by:
@@
where $H^{\prime}\times W^{\prime}<N$, with $K$, $r$ as small constants. Notably, for typical ViT configurations (e.g., ViT-B/16, $N=196$ and $d=768$), $N$ is considerably smaller than $d$. Consequently, VAPT can achieve greater parameter efficiency than VPT, as empirically shown in Table 1 and Table 2. Moreover, VAPT introduces only marginal computational overhead, up to 0.6% relative to VPT (see Appendix E.7).
A key strength of VAPT is that it achieves flexibility and effectiveness while maintaining computational efficiency. As discussed in Section 3, a naive linear prompt expert would require significantly more parameters due to the high dimensionality $Nd$ of the input $X$. In contrast, VAPT strikes an effective balance between parameter efficiency and prompt expert expressivity. Furthermore, while there is no principled design for adaptive prompts, VAPT’s architectural choices of token-wise projectors and channel-wise convolutions enable a straightforward implementation and simple prompt expert formulation. This inherent simplicity is pivotal, as it facilitates the thorough theoretical examination presented in Section 5, an aspect largely absent from prior prompt-based learning studies.
@@ 5 Statistical Advantages of VAPT
In this section, we explore the theoretical advantages of VAPT through its MoE connection, as detailed in Equation (7). This perspective provides a rigorous framework for analyzing the convergence properties of prompt estimation in MoE models le2024mixture ; le2024revisiting . Recalling from Section 3 that the MoE models $\tilde{{\bm{h}}}_{m,1},\dots,\tilde{{\bm{h}}}_{m,N}$ in each attention head share a common structure of experts and score functions. Furthermore, as noted in Section 2.1, omitting $\tilde{{\bm{h}}}_{m,N+1},\dots,\tilde{{\bm{h}}}_{m,N+N_{p}}$ does not affect ViT block output. Thus, to simplify our analysis while maintaining rigor, we focus on the first head (i.e., $m=1$), and specifically the first row of its attention matrix (i.e., $i=1$) in Equation (7). Within this simplified setting, we consider a regression framework for MoE models as follows.
@@
[…]
ViT-B/16 dosovitskiy2020image Tuned/ Scope Extra VTAB-1K zhai2019large [19]
Full iofinova2022well 100.00 88.54 75.88 83.36 47.64 65.57
Linear iofinova2022well 0.08 79.32 [0] 68.93 [1] 77.16 [1] 26.84 [0] 52.94
Partial-1 yosinski2014transferable 8.34 82.63 [0] 69.44 [2] 78.53 [0] 34.17 [0] 56.52
MLP-3 chen2020improved 1.44 79.80 [0] 67.80 [2] 72.83 [0] 30.62 [0] 53.21
Sidetune zhang2020side 10.08 78.35 [0] 58.21 [0] 68.12 [0] 23.41 [0] 45.65
Bias rebuffi2017learning 0.80 88.41 [3] 73.30 [3] 78.25 [0] 44.09 [2] 62.05
Adapter cai2020tinytl 1.02 85.46 [1] 70.67 [4] 77.80 [0] 33.09 [0] 62.41
LoRA hu2021lora 0.73 89.46 [3] 78.26 [5] 83.78 [2] 56.20 [7] 72.25
VPT-Shallow jia2022visual 0.16 84.62 [1] 76.81 [4] 79.66 [0] 46.98 [4] 64.85
VPT-Deep jia2022visual 0.73 89.11 [4] 78.48 [6] 82.43 [2] 54.98 [8] 69.43
E2VPT han20232vpt 0.39 89.22 [4] 80.01 [6] 84.43 [3] 57.39 [8] 71.42
VAPT (Ours) 0.36 89.58 [4] {4} 81.43 [6] {7} 85.13 [4] {4} 59.34 [8] {8} 72.91
Least-Squares Estimator: To estimate the unknown prompt parameters or, equivalently, the ground-truth mixing measure $G_{*}$, we use the least-squares method (vandeGeer-00, ). In particular, we consider the estimator defined as follows:
@@
where $\mathcal{G}_{L^{\prime}}(\Theta):=\{G=\sum_{i=1}^{\ell}\exp(b_{i})\delta_{(W_{1,i},W_{2})}:\ell\in[L^{\prime}],\ (b_{i},W_{1,i},W_{2})\in\Theta\}$ is the set of all mixing measures with at most $L^{\prime}$ atoms. Since the true number of experts $L$ is generally unknown, we assume the number of fitted experts $L^{\prime}$ is sufficiently large, i.e., $L^{\prime}>L$. To analyze prompt estimation convergence, it is essential to define a suitable loss function on the prompt parameters. In this work, we propose the Voronoi loss function, derived from the concept of Voronoi cells (manole22refined, ).
@@
$\mathcal{V}_{j}:=\left\{i\in[L^{\prime}]:\|Z_{i}-Z_{*,j}\|\leq\|Z_{i}-Z_{*,\ell}\|,\forall\ell\neq j\right\},$
@@ 6.1 Experimental Setup
Datasets. We evaluate our approach on two benchmarks: FGVC and VTAB-1K zhai2019large . FGVC includes five fine-grained classification datasets: CUB wah2011caltech , Oxford Flowers nilsback2008automated , Stanford Cars gebru2017fine , Stanford Dogs khosla2011novel , and NABirds van2015building , which challenge models to distinguish visually similar classes. VTAB-1K comprises 19 datasets, each with 1,000 training examples, grouped into three categories: Natural (images captured by standard cameras), Specialized (images collected via specialized equipment), and Structured (tasks requiring structural understanding, such as 3D depth prediction). Beyond classification, we also assess performance on semantic segmentation using ADE20K zhou2019semantic .
Baselines. In line with prior work jia2022visual ; han20232vpt , we compare VAPT against commonly used PEFT methods. All classification experiments use standard Vision Transformer (ViT) dosovitskiy2020image that are supervised pre-trained on ImageNet-21K deng2009imagenet . For semantic segmentation, we employ SETR-PUP zheng2021rethinking , which uses ViT as the backbone encoder. Additionally, we evaluate VAPT with backbones pre-trained using two self-supervised learning: MAE he2022masked and MoCo v3 chen2021empirical .
Training. We perform a grid search on val set of each task to determine the optimal learning rate, weight decay, kernel size $K$, and projector dimension $r$. We schedule the learning rate using a cosine decay schedule and train all models for 100 epochs. Following jia2022visual , we use batch sizes of 64 and 128. Additional implementation details can be found in Appendix D.
@@ 6.2 Empirical Results
[…]
Pre-trained objectives MAE he2022masked MoCo v3 chen2021empirical
Tuned/ VTAB-1K zhai2019large [19] Tuned/ VTAB-1K zhai2019large [19]
@@
Full iofinova2022well 100.00 59.31 79.68 53.82 100.00 71.95 84.72 51.98
Linear iofinova2022well 0.04 18.87 [0] 53.72 [0] 23.70 [0] 0.04 67.46 [4] 81.08 [0] 30.33 [0]
Partial-1 yosinski2014transferable 8.30 58.44 [5] 78.28 [1] 47.64 [1] 8.30 72.31 [5] 84.58 [2] 47.89 [1]
Bias rebuffi2017learning 0.16 54.55 [1] 75.68 [1] 47.70 [0] 0.16 72.89 [3] 81.14 [0] 53.43 [4]
Adapter cai2020tinytl 0.87 54.90 [3] 75.19 [1] 38.98 [0] 1.12 74.19 [4] 82.66 [1] 47.69 [2]
VPT-Shallow jia2022visual 0.05 39.96 [1] 69.65 [0] 27.50 [0] 0.06 67.34 [3] 82.26 [0] 37.55 [0]
VPT-Deep jia2022visual 0.31 36.02 [0] 60.61 [1] 26.57 [0] 0.22 70.27 [4] 83.04 [0] 42.38 [0]
GateVPT yoo2023improving 0.05 47.61 [2] 76.86 [1] 36.80 [1] 0.06 74.84 [4] 83.38 [1] 49.10 [3]
@@
Overall Comparison. Table 1 compares VAPT’s performance against full fine-tuning and other prominent PEFT methods on VTAB-1K and FGVC. Full iofinova2022well denotes full fine-tuning, which updates all model parameters, while methods such as Linear iofinova2022well , Partial-1 yosinski2014transferable (top layer), and MLP-3 chen2020improved (3 MLP layers) modify only a subset of parameters. Sidetune zhang2020side , Bias rebuffi2017learning , Adapter cai2020tinytl , and LoRA hu2021lora introduce trainable modules for adaptation. Concurrent visual prompt tuning approaches, VPT jia2022visual and E2VPT han20232vpt , are also included for comparison (see Appendix C for further details). Our results indicate that VAPT surpasses full fine-tuning in 22 out of 24 tasks. Specifically, VAPT achieves a notable 1.04% accuracy increase on FGVC and a substantial 11.70% improvement on VTAB-1K Structured. Across the entire VTAB-1K benchmark, VAPT demonstrates an average gain of 7.34% over full fine-tuning, while updating merely 0.36% of the backbone parameters. These findings highlight VAPT’s significant effectiveness and efficiency as an innovative PEFT method. Additionally, VAPT achieves state-of-the-art performance compared to other PEFT approaches. Among prompt tuning methods, VAPT consistently outperforms VPT in 23 out of 24 tasks, despite utilizing fewer parameters. We attribute these improvements to VAPT’s effective design, which enhances the expressiveness of prompt experts. These results underscore VAPT’s potential as a powerful tool for improving model performance with significantly reduced parameter overhead.
Different Pre-training Methods. We investigate VAPT’s performance when initialized with backbones pre-trained using different self-supervised learning (SSL) objectives, specifically MAE he2022masked and MoCo v3 chen2021empirical , with detailed results in Table 2. Previous studies jia2022visual ; yoo2023improving have indicated that VPT can exhibit suboptimal performance with SSL pre-trained backbones. In contrast, VAPT achieves significant performance improvements by effectively leveraging the rich information present in the input features. For example, VAPT demonstrates a remarkable 23.21% accuracy improvement under MAE on VTAB-1K Natural tasks and an 18.36% improvement with MoCo v3 on VTAB-1K Structured. Furthermore, when compared to other PEFT methods, VAPT consistently outperforms them, achieving the highest “Number of Wins” relative to full fine-tuning, with 9 out of 19 tasks under MAE and 15 out of 19 tasks under MoCo v3. These results underscore the generality and robustness of our method across different pre-training objectives, supported by both theoretical analysis (Section 5) and empirical evidence.
@@
Table 3: Classification accuracy on Stanford Dogs khosla2011novel using varying fractions of training data. Bold indicates the best performance.
@@
Sample Efficiency. As theoretically established in Section 5, our proposed formulation exhibits optimal sample efficiency for prompt estimation. To empirically validate this, we systematic experiments on the Stanford Dogs dataset khosla2011novel . Following d2021convit , we subsampled each class by fraction $f=\{0.01,0.1,0.3,0.5,1.0\}$ and scaled the number of training epochs by $1/f$, thereby ensuring a constant total number of image presentations to the model. The results, presented in Figure 2 and Table 3, demonstrate that VAPT consistently outperforms VPT across all evaluated training set sizes. Notably, with only 1% of the data, VPT achieves an accuracy of 3.6%, whereas VAPT attains a substantially higher accuracy of 60.1%. Furthermore, VAPT requires only 30% of the data to match the performance of VPT trained on the full 100% dataset, signifying an approximate 3×3$\times$3 × reduction in data requirements. This underscores the superior sample efficiency of our approach.

### v6

@@ 4.2 Adaptive Prompt
Efficiency Considerations. Following VPT, we insert adaptive prompts into every ViT block. While the feature projector is a lightweight MLP, distinct projector for each block would incur substantial parameter and memory overhead. To mitigate this, we employ a single, shared feature projector $g(\cdot)$ reused across all ViT blocks. This significantly reduces the overall computational cost and enhances efficiency, while preserving prompt expert expressiveness relative to VPT. Furthermore, an independent LayerNorm (Lei Ba et al., 2016) is incorporated, we incorporate an independent LayerNorm (Lei Ba et al., 2016) before the token-wise projectors in each block. Regarding parameter complexity, VPT introduces approximately $L\times N_{p}\times d$ parameters across $L$ ViT blocks. In contrast, VAPT’s learnable parameters include token-wise projectors, a channel-wise convolution layer per block, and a shared feature projector. The total parameter count is given by:
@@
where $H^{\prime}\times W^{\prime}<N$, with $K$, $r$ as small constants. Notably, for typical ViT configurations (e.g., ViT-B/16, $N=196$ and $d=768$), NN is considerably smaller than $d$. Consequently, VAPT can achieve greater parameter efficiency than VPT, as empirically shown in Table 1 and Table 2. Moreover, VAPT introduces only marginal computational overhead, up to 0.6% relative to VPT (see Appendix E.7).
Novelty. A central contribution of VAPT is its ability to achieve both flexibility and effectiveness while preserving computational efficiency. A common assumption is that increasing functional expressiveness necessarily incurs substantial computational or parameter overhead. As outlined in the above analysis, a naive implementation of a linear prompt expert would require a prohibitively large number of parameters due to the high dimensionality ($Nd$) of the input $\bm{X}$}. In contrast, VAPT strikes a favorable balance between parameter efficiency and the expressivity of prompt experts. Despite the absence of a principled framework for designing adaptive prompts, VAPT’s architectural choices of token-wise projectors and channel-wise convolutions yield a straightforward implementation and a compact, interpretable prompt-expert formulation. This architectural simplicity is particularly important, as it enables the rigorous theoretical analysis in Section 5, where we show that VAPT achieves an optimal sample-efficiency rate. Prior work often relies on more complex architectures or heuristic mechanisms to adapt prompts across inputs (Wang et al., 2022; Huang et al., 2023; Kim et al., 2023). While these designs can perform well empirically, their functional complexity typically precludes meaningful theoretical understanding. By contrast, VAPT is deliberately designed to retain a simple yet expressive functional form, leading to the clean expert formulations in Equations (13) and (14). These formulations support a thorough theoretical analysis without sacrificing empirical performance, addressing a key gap in the literature: VAPT not only performs well in practice but also enjoys provable robustness and generalization guarantees. Importantly, our approach leverages the existing MoE structure implicit in attention heads rather than imposing an additional, external MoE module, as is common in prior work (Zeng et al., 2025). Conventional MoE-based methods explicitly insert new routing components and expert modules into the Transformer architecture, introducing substantial architectural modifications. In contrast, VAPT avoids such intrusive changes to the backbone while still admitting a formulation that is amenable to formal analysis. This perspective allows us to design VAPT as a simple and practical mechanism that fully benefits from MoE theory to rigorously characterize its behavior. For a more detailed comparison with related methods, please refer to Appendix C.
@@ 5 Statistical Advantages of VAPT
In this section, we explore the theoretical advantages of VAPT through its MoE connection, as detailed in Equation (7). This perspective provides a rigorous framework for analyzing the convergence properties of prompt estimation in MoE models (Le et al., 2024; 2025b). Recalling from Section 3 that the MoE models $\tilde{{\bm{h}}}_{m,1},\dots,\tilde{{\bm{h}}}_{m,N}$ in each attention head share a common structure of experts and score functions. Furthermore, as noted in Section 2.1, omitting $\tilde{{\bm{h}}}_{m,N+1},\dots,\tilde{{\bm{h}}}_{m,N+N_{p}}$ does not affect ViT block output. Thus, to simplify our analysis while maintaining rigor, we focus on the first head (i.e., $m=1$), and specifically the first row of its attention matrix (i.e., $i=1$) in Equation (7). Within this simplified setting, we consider a regression framework for MoE models as follows.
@@
Least-Squares Estimator: To estimate the unknown prompt parameters or, equivalently, the ground-truth mixing measure $G_{*}$, we use the least-squares method (van de Geer, 2000). In particular, we consider the estimator defined as follows:
@@
where $\mathcal{G}_{L^{\prime}}(\Theta):=\{G=\sum_{i=1}^{\ell}\exp(b_{i})\delta_{(W_{1,i},W_{2})}:\ell\in[L^{\prime}],\ (b_{i},W_{1,i},W_{2})\in\Theta\}$ is the set of all mixing measures with at most $L^{\prime}$ atoms. Since the true number of experts $L$ is generally unknown, we assume the number of fitted experts $L^{\prime}$ is sufficiently large, i.e., $L^{\prime}>L$. To analyze prompt estimation convergence, it is essential to define a suitable loss function on the prompt parameters. In this work, we propose the Voronoi loss function, derived from the concept of Voronoi cells (Manole & Ho, 2022).
@@
$\hskip-5.0pt\mathcal{V}_{j}:=\left\{i\in[L^{\prime}]:\|Z_{i}-Z_{*,j}\|\leq\|Z_{i}-Z_{*,\ell}\|,\forall\ell\neq j\right\},$
@@
Table 1: Overall Comparison for ViT-B/16 Supervised Pre-trained on ImageNet-21K. Following Jia et al. (2022), we report average accuracy (3 runs) on FGVC and VTAB-1K, “Number of Wins” $\cdot$] compared to full fine-tuning, and “Number of Wins over VPT” $\left\{\cdot\right\}$. “Tuned/Total” denotes the average percentage of parameters tuned (24 tasks), “Scope” specifies the tuning scope, and “Additional parameters” indicates if parameters beyond pre-trained backbone/head are introduced. Per-task results are in Appendix E.1.
ViT-B/16 (Dosovitskiy, 2020) Tuned/ Scope Extra VTAB-1K (Zhai et al., 2019) [19]
Full (Iofinova et al., 2022) 100.00 88.54 75.88 83.36 47.64 65.57
Linear (Iofinova et al., 2022) 0.08 79.32 [0] 68.93 [1] 77.16 [1] 26.84 [0] 52.94
Partial-1 (Yosinski et al., 2014) 8.34 82.63 [0] 69.44 [2] 78.53 [0] 34.17 [0] 56.52
MLP-3 (Chen et al., 2020) 1.44 79.80 [0] 67.80 [2] 72.83 [0] 30.62 [0] 53.21
Sidetune (Zhang et al., 2020) 10.08 78.35 [0] 58.21 [0] 68.12 [0] 23.41 [0] 45.65
Bias (Rebuffi et al., 2017) 0.80 88.41 [3] 73.30 [3] 78.25 [0] 44.09 [2] 62.05
Adapter (Cai et al., 2020) 1.02 85.46 [1] 70.67 [4] 77.80 [0] 33.09 [0] 62.41
LoRA (Hu et al., 2021) 0.73 89.46 [3] 78.26 [5] 83.78 [2] 56.20 [7] 72.25
VPT-Shallow (Jia et al., 2022) 0.16 84.62 [1] 76.81 [4] 79.66 [0] 46.98 [4] 64.85
VPT-Deep (Jia et al., 2022) 0.73 89.11 [4] 78.48 [6] 82.43 [2] 54.98 [8] 69.43
E2VPT (Han et al., 2023) 0.39 89.22 [4] 80.01 [6] 84.43 [3] 57.39 [8] 71.42
SA2VP (Pei et al., 2024) 0.65 90.08 [4] 80.97 [6] 85.73 [4] 60.80 [8] 73.47
ViaPT (Xiao et al., 2025) 0.66 91.40 [4] 82.62 [6] 85.22 [2] 61.25 [8] 73.70
VFPT (Zeng et al., 2024) 0.66 89.24 [4] 81.35 [6] 84.93 [4] 60.19 [8] 73.20
VAPT (Ours) 0.36 89.58 [4] {4} 81.43 [6] {7} 85.13 [4] {4} 59.34 [8] {8} 72.91
@@ 6.1 Experimental Setup
Datasets. We evaluate VAPT on two benchmarks: FGVC and VTAB-1K (Zhai et al., 2019). FGVC includes five fine-grained classification datasets: CUB (Wah et al., 2011), Oxford Flowers (Nilsback & Zisserman, 2008), Stanford Cars (Gebru et al., 2017), Stanford Dogs (Khosla et al., 2011), and NABirds (Van Horn et al., 2015) that require distinguishing between visually similar classes. VTAB-1K comprises 19 datasets, each with 1,000 training examples, grouped into: Natural (images captured by standard cameras), Specialized (images collected via specialized equipment), and Structured (tasks requiring structural understanding, such as 3D depth prediction). Beyond classification, we also assess performance on semantic segmentation using ADE20K (Zhou et al., 2019).
Baselines. In line with prior work (Jia et al., 2022; Han et al., 2023), we compare VAPT against commonly used PEFT methods. All classification experiments use standard Vision Transformer (ViT) (Dosovitskiy, 2020) that are supervised pre-trained on ImageNet-21K (Deng et al., 2009). For semantic segmentation, we employ SETR-PUP (Zheng et al., 2021), which uses ViT as the backbone encoder. Additionally, we evaluate VAPT with backbones pre-trained using two self-supervised learning: MAE (He et al., 2022) and MoCo v3 (Chen et al., 2021).
Training. We perform a grid search on val set of each task to determine the optimal learning rate, weight decay, kernel size $K$, and projector dimension $r$. We schedule the learning rate using a cosine decay schedule and train all models for 100 epochs. Following Jia et al. (2022), we use batch sizes of 64 and 128. Additional implementation details can be found in Appendix D.
@@ 6.2 Empirical Results
Table 2: Comparison of Different Pre-training Objectives. We consider MAE (He et al., 2022) and MoCo v3 (Chen et al., 2021) using ViT-B/16. We report test accuracy on VTAB-1K, the “Number of Wins” $\cdot$] relative to full fine-tuning and the “Number of Wins over VPT” $\left\{\cdot\right\}$. “Tuned/Total” denotes the average percentage of parameters tuned. Per-task results are in Appendix E.2.
Pre-trained objectives MAE (He et al., 2022) MoCo v3 (Chen et al., 2021)
Tuned/ VTAB-1K (Zhai et al., 2019) [19] Tuned/ VTAB-1K (Zhai et al., 2019) [19]
@@
Full (Iofinova et al., 2022) 100.00 59.31 79.68 53.82 100.00 71.95 84.72 51.98
Linear (Iofinova et al., 2022) 0.04 18.87 [0] 53.72 [0] 23.70 [0] 0.04 67.46 [4] 81.08 [0] 30.33 [0]
Partial-1 (Yosinski et al., 2014) 8.30 58.44 [5] 78.28 [1] 47.64 [1] 8.30 72.31 [5] 84.58 [2] 47.89 [1]
Bias (Rebuffi et al., 2017) 0.16 54.55 [1] 75.68 [1] 47.70 [0] 0.16 72.89 [3] 81.14 [0] 53.43 [4]
Adapter (Cai et al., 2020) 0.87 54.90 [3] 75.19 [1] 38.98 [0] 1.12 74.19 [4] 82.66 [1] 47.69 [2]
VPT-Shallow (Jia et al., 2022) 0.05 39.96 [1] 69.65 [0] 27.50 [0] 0.06 67.34 [3] 82.26 [0] 37.55 [0]
VPT-Deep (Jia et al., 2022) 0.31 36.02 [0] 60.61 [1] 26.57 [0] 0.22 70.27 [4] 83.04 [0] 42.38 [0]
GateVPT (Yoo et al., 2023) 0.05 47.61 [2] 76.86 [1] 36.80 [1] 0.06 74.84 [4] 83.38 [1] 49.10 [3]
E2VPT (Han et al., 2023) 0.07 59.52 [4] 77.80 [1] 44.65 [3] 0.13 76.47 [4] 87.28 [2] 54.91 [6]
ViaPT (Xiao et al., 2025) 0.36 54.26 [-] 78.01 [-] 37.52 [-] 0.30 79.12 [-] 86.81 [-] 60.05 [-]
VFPT (Zeng et al., 2024) 0.38 53.59 [6] 77.75 [1] 36.15 [1] 0.22 77.47 [5] 85.76 [3] 58.74 [6]
@@
Overall Comparison. Table 1 compares VAPT against full fine-tuning and other prominent PEFT methods on VTAB-1K and FGVC. Full denotes full fine-tuning, which updates all model parameters, while methods such as Linear, Partial-1 (top layer), and MLP-3 (3 MLP layers) modify only a subset of parameters. Sidetune (Zhang et al., 2020), Bias (Rebuffi et al., 2017), Adapter (Cai et al., 2020), and LoRA (Hu et al., 2021) introduce trainable modules for adaptation. Concurrent visual prompt tuning approaches, VPT (Jia et al., 2022), E2VPT (Han et al., 2023), SA2VP (Pei et al., 2024), ViaPT (Xiao et al., 2025) and VFPT (Zeng et al., 2024), are also included for comparison (see Appendix C for further details). Our results indicate that VAPT surpasses full fine-tuning in 22 out of 24 tasks. Specifically, VAPT achieves a notable 1.04% accuracy increase on FGVC and a substantial 11.70% improvement on VTAB-1K Structured. Across the entire VTAB-1K benchmark, VAPT demonstrates an average gain of 7.34% over full fine-tuning, while updating merely 0.36% of the backbone parameters. These findings highlight VAPT’s significant effectiveness and efficiency as an innovative PEFT method. Additionally, VAPT achieves state-of-the-art performance compared to other PEFT approaches. Among prompt tuning methods, VAPT consistently outperforms VPT in 23 out of 24 tasks and attains competitive performance with VFPT, a recent state-of-the-art approach, despite utilizing nearly 50% fewer parameters. We attribute these improvements to VAPT’s design, which enhances the expressiveness of prompt experts. These results underscore VAPT’s potential as a powerful tool for improving performance with significantly reduced parameter overhead.
Different Pre-training Methods. We investigate VAPT’s performance when initialized with backbones pre-trained using different self-supervised learning (SSL) objectives, specifically MAE (He et al., 2022) and MoCo v3 (Chen et al., 2021), with detailed results in Table 2. Previous studies (Jia et al., 2022; Yoo et al., 2023) have indicated that VPT can exhibit suboptimal performance with SSL pre-trained backbones. In contrast, VAPT achieves significant performance improvements by effectively leveraging the rich information present in the input features. For example, VAPT demonstrates a remarkable 23.21% accuracy improvement under MAE on VTAB-1K Natural tasks and an 18.36% improvement with MoCo v3 on VTAB-1K Structured. Furthermore, when compared to other PEFT methods, VAPT consistently outperforms them, achieving the highest “Number of Wins” relative to full fine-tuning, with 9 out of 19 tasks under MAE and 15 out of 19 tasks under MoCo v3. In addition, VAPT outperforms state-of-the-art approaches such as VFPT under MAE and remains competitive under MoCo v3. These results underscore the generality and robustness of our method across different pre-training objectives, supported by both theoretical analysis (Section 5) and empirical evidence.
@@
Table 3: Classification accuracy on Stanford Dogs (Khosla et al., 2011) using varying fractions of training data. Bold indicates the best performance.
@@
Sample Efficiency. To empirically validate the theoretical sample efficiency from Section 5, we conducted experiments on the Stanford Dogs dataset (Khosla et al., 2011). Following d’Ascoli et al. (2021), we subsampled each class by fraction $f=\{0.01,0.1,0.3,0.5,1.0\}$ and scaled the number of training epochs by $1/f$, thereby ensuring a constant total number of image presentations to the model. The results, presented in Figure 2 and Table 3, demonstrate that VAPT consistently outperforms VPT across all evaluated training set sizes. Notably, with only 1% of the data, VPT achieves an accuracy of 3.6%, whereas VAPT attains a substantially higher accuracy of 60.1%. Furthermore, VAPT requires only 30% of the data to match the performance of VPT trained on the full 100% dataset, signifying an approximate $3\times$ reduction in data requirements. This underscores the superior sample efficiency of our approach.
It is important to note that the primary objective of this work is not to establish a new state-of-the-art on every dataset. As outlined in the abstract and introduction, our focus is on scholarly value rather than purely incremental methodological gains: we aim to provide the community with clear insights into the benefits of increasing the functional expressiveness of prompt experts within visual prompt tuning frameworks. VAPT serves as a concrete instantiation of this idea. Its adaptive formulation illustrates the potential of adaptive prompt experts to deliver improved performance, stronger parameter efficiency, and enhanced sample efficiency in both theory and practice.
Within this perspective, the most relevant comparison is between VAPT’s adaptive prompt-expert formulation and VPT, where prompt experts are constant functions. Our experiments consistently show that VAPT outperforms VPT on the majority of benchmarks while using a more compact set of trainable parameters. Moreover, our empirical findings quantify the statistical advantages of VAPT over VPT, and these observations are rigorously supported by our theoretical analysis in Section 5 and Appendix A. Together, these results reinforce our central claim that increasing the expressiveness of prompt experts is a principled and effective direction for advancing visual prompt tuning.